thehearken:

anyone have tips on how to remember how to solve for volume and surface area in prisms and cylinders? I'm having trouble understanding and would like to know if anyone has any easier to understand methods

thehearken:

(I'm in geometry, just for anyone who might wanna reply lol...don't want replies for other courses I'm not in yet..)

EdwinJsHispanic:

for cylinders I'm pretty sure its v= pie(R) to the power of 2 times H

thehearken:

@edwinjshispanic wrote:
for cylinders I'm pretty sure its v= pie(R) to the power of 2 times H
well yes, I figured that out... I meant is there any easier ways to remember the equations?

EdwinJsHispanic:

uhm. No I don't think so, I think js remembering the formula is it sorry

thehearken:

darn.. that's fine. I'm just having trouble understanding is all, and my teacher doesn't explain it that well haha. thanks though

EdwinJsHispanic:

lol ofc

Twaylor:

for finding the volume of a cylinder its V=πr2h, for total surface area it's A=2πrh+2πr2, the total surface area of any prism you can think of is the sum of the areas of all its outer faces. the universal formula is: SA = 2B +Ph (SA being the surface area, B being the area of one base, P being the perimeter of one base, h being the height (length) of said prism) the volume of any prism is V = B*H (the volume of any prism is calculated by multiplying the area of its base by its height, the base area depends on the shape of the prism's cross-section, while the is the distance between the two parallel bases):3

thehearken:

thank you twaylor so helpful and amazing

Twaylor:

@thehearken wrote:
thank you twaylor so helpful and amazing
yuippie :3

EdwinJsHispanic:

damn, smartie pants

DUXK:

@edwinjshispanic wrote:
damn, smartie pants
language brochacho

EdwinJsHispanic:

this brochacho has a name boi

thehearken:

@duxk wrote:
@edwinjshispanic wrote:
damn, smartie pants
language brochacho
damn isnt a bad word if it was it woould be censored by the coding of the site

Twaylor:

@thehearken wrote:
@duxk wrote:
@edwinjshispanic wrote:
damn, smartie pants
language brochacho
damn isnt a bad word if it was it woould be censored by the coding of the site
there are plenty of slurs the moderators missed that even experienced sailors wouldn't say.

DUXK:

@edwinjshispanic wrote:
this brochacho has a name boi
ik n don't "boi" me

EdwinJsHispanic:

I can boi u all I want